OpenStudy (mackenzie2013):

What is the input value for the following function if the output value is 4.5? y = x/2 2.25 6.5 8 9

OpenStudy (anonymous):

output is y input is x

OpenStudy (anonymous):

The input value would be the x, and the output value would be the y. So, given y=x/2, we need to find x if y=4.5. So we'll plug in 4.5 for y: \[4.5=x/2\] So we only need to multiply both sides by 2 to get our answer: \[9=x\]

OpenStudy (cwrw238):

solve for x 4. 5 = x/2